OpenStudy (anonymous):

A painting is 6 cm longer than it is wide. The painting is to be surrounded by a mat that is 2 cm wide and covered by a piece of glass with an area of 352cm squared. Find the dimensions of the painting. (A verbal model and equation is required)

OpenStudy (anonymous):

L = W + 6 (1st sentence) A = L * W With the mat L = W + 10 (2 cm more on each side) W = W + 4 (2 cm more on each side) Area of mat - A = L* W A = (W + 10)(W + 4) A = W^2 + 4W + 10W + 40 A = W^2 + 14W + 40 352 = W^2 + 14W + 40 0 = W^2 + 14W - 312 0 = (W + 26)(W - 12) W = -26 or W = 12 Can't use a negative width W = 12 L = 18 (from first sentence L = W + 6)